OpenStudy (anonymous):

divide and write in scientific notation 6 *10^-4/3*10^-6

OpenStudy (agreene):

\[\frac{6*10^{-4}}{3*10^{-6}}=\frac{0.0006}{0.000003}=200\] \[=2*10^2\]

OpenStudy (anonymous):

thanks